Q 1: A cube of soil measures 0.1x0.1x0.1m and has a wet mass of 1.46 kg, of which 0.26 kg is water. Find: (a) gravimetric and (b) volumetric water content, (c) equivalent depth of water, (c) bulk density, (d) total porosity, and (e) relative saturation. (Note: assume data not provided)
Answer #1

A) Gravimetric water content = weight of water/weight of solid soil

= 0.26kg/(1.46-0.26)kg = 0.2168 or 21.67%

b) we know density of dry soil = 2650 kg/m3

volumetric content of water = (weght of water/density of water )/ (weght of dry soil / density of dry soil )

or volume of water/volume of dry soil

= (0.26 kg/1000kg/m3)(1.20kg/2650kg/m3)

= 0.26*2.65/1.20 = 0.5742 or 57.41%

c) bulk density = weight of soil / total volume

= 1.46 kg/ 0.001m3 = 1460 kg/m3

d) total   porosity = volume of void / total volume

= (0.001-1.20/2650)/0.001

= 0.5472 or 54.72%

e) relative saturation = volume of water / volume of void

=(0.26kg/1000kg/m3)/(0.001-1.20/2650kg/m3)

= 0.4751 or 47.51%